# Reconfigurable Intelligent Surfaces for 6G Mobile Networks: An Industry R&D Perspective

### v2

@@ II-C Technical Challenges
Despite the considerable potential of RISs in revolutionizing next-generation networks, several challenges need to be addressed, which might prevent a successful integration into mobile networks. This section provides a comprehensive discussion of these challenges and refers to recent literature.
@@ II-C1 Channel Estimation and Radio Signal Overhead
Accurate channel state information (CSI) is vital for the efficient operation of RISs. This information enables the control of passive elements to achieve optimal signal reflection and manipulation. As narrow beams are required to overcome path loss, a high number of RIS elements is needed. This leads to a high control signaling overhead for the controlling of the RIS. The already challenging overhead is further affected by the following effects: In general, a cascaded channel between the base station, RIS, and UE must be estimated. However, obtaining accurate CSI of this channel is challenging due to potential near-field characteristics, mutual coupling, and beam squint effects, which can lead to errors in channel estimation and performance degradation. Furthermore, recently published study on optimal RIS configuration in a rich scattering environment indicates, that the linear cascaded models is not correct [18, 19]. All named challenges add up in increasing the required overhead to optimal configure the RIS. To address the challenge of a tremendous overhead, passive beam training techniques employing predefined beam pattern codebooks were suggested [20, 21, 22, 23]. These approaches seek to exploit the inherent structure of the RIS-assisted channel to obtain accurate CSI required for optimal RIS manipulation but misses the ability for flexible deployments. Further research must investigate for balancing the trade-off between RIS coverage and resource overhead while maintaining communication quality.
@@ II-C2 Regulatory and Multi-Operator Interference Issues
One of the major challenges for RIS deployment in multi-operator scenarios is managing interference among different network operators, ensuring fairness in resource allocation, and avoiding destructive interference. As RISs can potentially modify the propagation environment across multiple operators, the allocation of resources and the concurrent usage of multiple RISs deployed by different operators become critical concerns. The wideband properties and non-linear saturation effects of RISs introduce challenges related to unwanted interference and regulatory compliance [20]. When an RIS is deployed and controlled by one operator, it can cause unpredictable changes in the propagation environment for other operators in the nearby spectrum. Such alterations in the propagation environment can lead to issues such as additional reflections affecting cell coverage, neighbour-cell interference, and may disturb network planning. Furthermore, rapid variations in reflection patterns can disturb instantaneous spatial multiplexing and link adaptation for both reciprocity-based and downlink CSI-based scenarios. Research in this area is in its nascent stages, but preliminary work suggests the need for collaborative efforts among operators. Addressing these challenges necessitates the development of interference mitigation techniques and careful consideration of regulatory guidelines.
Possible solutions for mitigating this issue include the development of standardized protocols for RIS deployment and interference management, enhancement of cooperation schemes like cooperative scheduling among network operators, and interference-aware resource allocation algorithms that minimize multi-operator interference. Further research must focus on identifying and quantifying the key factors influencing multi-operator interference and exploring advanced techniques for interference mitigation and management in RIS-assisted networks.
@@ II-C3 Competitive Technologies
The benefits of RISs must also be assessed compared to competitive technologies like the decode-and-forward relay [24, 25] or the network-controlled repeater (NCRs) [26]. While these solutions may be more costly alternatives compared to RIS, they offer better controllability and potentially larger coverage, as RISs are more susceptible to multiplicative fading. The path loss of the BS-RIS/NCR-UE link is a product of the individual link path losses. It can be compensated by amplification capabilities in the NCR case. For the case of RIS, it requires a close deployment of the RIS to either the BS or the UE to enable a signal improvement or a very large RIS with an extensive element count. Besides the production costs, deployment issues and maintenance costs, which all increase with size, a very large RIS leads to problems described in the earlier paragraph on channel estimation challenges. For a fair comparison between RIS and NCRs all named advantages and disadvantages must be added up to a total cost of ownership per deployment. To evaluate the economic choice between RIS and NCR, the network operators will have to consider the cost for coverage, which is a sum of the deployment costs per area. Because several cheaper RISs are required to cover the same area as one NCR, as stated earlier, the lower RIS deployment cost must level out this factor. The many aspects of the cost for coverage lead to a complex calculation. This evaluation is seen as a crucial aspect of future research. A more extensive comparison of RISs and NCRs is provided in [7].
@@ III Hardware Design
RISs are based on unit cell elements that apply a phase shift to the incoming waves and reradiate them back to the cell environment. In this section, hardware concepts that are under development in the 6G LICRIS project are briefly discussed to illustrate novel approaches to RIS design. In particular, the proposed concepts follow a passive design, reflecting the outgoing RF wave and saving power compared to base station antennas, with the goal of enhancing coverage, spectral efficiency, and beam management while improving energy efficiency. Two different RIS models are proposed at the 28 GHz band, which could be tested in indoor scenarios.
(a) Front view of RIS.
(b) Phase of the reflected wave for the unit cell between 22 and 35 GHz when the liquid crystal has different values of relative permittivity.
Figure 3: Phase reflection performance multi-resonant RIS design.
The first model consists of a multi-resonant RIS integrated in liquid crystal. This topology consists of a layer of liquid crystal comprised between two layers of glass substrate. Different metallic resonant structures are printed on the glass making use of photolitography. This way different resonant structures are on top of and below the liquid crystal layer, creating a multi-resonant topology whose resonant frequencies depend on the polarization of the molecules of the LC-material. When multiple resonances are used, a wide bandwidth is expected, in this case a bandwidth of around 6 GHz with a central frequency of 28 GHz is expected. This large bandwidth will enable the use of a high data-rate. This topology is compatible with a single polarization. A frontview of how the RIS would look like and the expected phase of a single unit cell of the RIS topology are shown in Figure 3. The phase different works between 24 and 32 GHz approximately with a difference of around 300, which is relatively constant in the whole frequency range.
(a) Stackup view of a unit cell of the stacked-patch RIS
(b) Reflected phase for the unit cell based on reconfigurable phase shifters between 22 and 35 GHz when the liquid crystal has different values of relative permittivity
Figure 4: Phase reflection performance of coupled-patch RIS design.
The second type of RIS consists of a coupled patch RIS. In this topology the radiating element is a patch which is coupled to a reflective LC phaseshifter. Once the wave reaches the patch, it is coupled into the active phaseshifter through a slot in the ground plane. This phaseshifter provides a changeable phaseshift which is between 0 and 360. A bandwidth of around 3 GHz is expected and it is compatible with two orthogonal polarizations. The whole stack-up of the RIS and the phaseshifter for different values of permittivity are shown in Figure 4.
@@ IV Deployment and Dimensioning Aspects
This section discusses the challenges related to RIS deployment and dimensioning. One of the main challenges in RIS deployment is related to calculation of RIS placement in a cellular network, while RIS dimensioning involves the calculation of the number of reflecting elements necessary to achieve the required system targets. Central to deployment and dimensioning is the propagation model considered in the calculations. To this end, this section begins with a discussion on RIS propagation modeling, followed by deployment optimization, and RIS dimensioning.
@@ IV-A Propagation Modeling
In order to manage and control RISs from a system-level perspective, it is necessary to understand and ideally anticipate the effects of their configuration on intended changes in radio propagation as well as unintended effects, e.g., on competing networks or other frequency bands. This requires fast and reliable RF propagation models, where a reasonable loss of accuracy may be acceptable.
Semi-empirical and/or geometric RF propagation models offer a good compromise. The measures that need to be taken into account are first of all the received power (see Figure 5), in a more sophisticated model, beyond the scope of the current work, also the impact of the MIMO path through the RIS are of interest. Since RIS can be most efficiently used in higher frequency beyond 6 GHz, RF propagation in scenarios where RIS are used is typically limited by line-of-sight (LoS) between transmitter and receiver. It is the intention of using RIS to supplement the LoS-pathes by reflections on RIS. While semi-empirical models are widely used in sub-6GHz-mobile radio scenarios (or historically sub-2GHz), geometric models are less limited in its frequency application ranges, since they are based on deterministic approaches.
Figure 5: Path model of RF propagation components including an RIS working as a “directed mirror”. The blocks named with “L” meaning losses, “G” gains and “P” is representing power values. Tx is transmitting, Rx is receiving. The loss of the RIS depends on the angle of arriving/transmitting of sender/receiver.
Semi-empirical models offer the advantage of relatively low computational complexity compared to other RF propagation modeling approaches such as ray tracing and full-wave simulations, which can be computationally intensive and time-consuming, and also have limitations. They are based on empirical measurements that may not always capture the full complexity of real-world environments, and they may not accurately predict the effects of new technologies and deployment scenscenarios that differ from the assumptions of the original model. Therefore, careful validation and calibration of semi-empirical models using real-world measurements and data is essential to ensure their accuracy and reliability.
@@
Geometric models are based on the concept that RF signals propagate in straight lines and their behavior can be predicted by considering the geometry of the environment in which they propagate (see Figure 6, right-hand side). These models use parameters such as the distance between the transmitter and receiver, the height of the antennas, and the surrounding terrain to calculate the path loss of RF signals. For example, the two-beam ground reflection model accounts for reflections of RF signals from the ground by assuming that the signal travels in two paths: a direct path from the transmitter to the receiver and another reflected path from the ground to the receiver. The model takes into account the heights of the antennas and the distance between the transmitter and receiver to calculate the path loss. The two-beam ground reflection model is commonly used to predict RF propagation in urban environments where ground reflections can significantly affect the path loss. A similar approach can be considered for RF prediction with RIS, assuming that the reflection point is from the RIS rather than the ground.
@@
More advanced geometric models can also be considered that include additional factors such as the height and location of buildings, the density of obstructions and the surrounding terrain. The Extended Hata Model [27] combines geometric principles with empirical data to account for the effects of both urban and suburban environments, including factors such as building density and road width. It can also account for the effects of obstacles, as is typically done in knife-edge models.
Knife-edge models are named for the sharp edge of the obstacle, which is considered to be the most significant factor affecting the diffraction of RF signals (see Figure 6, left-hand side). The shape, height and distance of the obstacle from the transmitter and receiver are key parameters in these models. One of the most commonly used knife-edge models is the single knife-edge diffraction model, which assumes that the RF signal encounters a single sharp edge along its propagation path. The model calculates the path loss based on the height and distance of the obstacle, the height of the antennas, the frequency of the RF signal, and the wavelength of the signal.
For the integration of RIS into RF propagation models, diffraction is certainly of less interest than reflection. However, both modeling approaches could be combined by a combined parameter estimation of their parameters with methods of mathematical optimization, including machine learning/artificial intelligence approaches (see Figure 7). Such an approach depends strongly on the availability of suitable training data. For RIS such data are currently hardly available. A combination of simulation and ray tracing may provide artificial training data, similar to the approach of RadioUNet [28].
In any case, the RF characteristics of RIS need to be well understood for different bands, in different environments and also under different environmental conditions. The proposed RF models need to be flexible in this respect, which drives the need for simplified models. For control approaches, RF measurement data from the field must constantly feed the control loops, and the RF modeling on the RIS must also be adapted during operations. Therefore, a complete control loop and measurement collection approach from connected devices within the network is required.

### v4

@@ II-C Technical Challenges
Despite the considerable potential of RISs in revolutionizing next-generation networks, several challenges need to be addressed, which might prevent successful integration into mobile networks. This section provides a comprehensive discussion of these challenges and refers to recent literature.
@@ II-C1 Channel Estimation and Radio Signal Overhead
Accurate channel state information (CSI) is vital for the efficient operation of RISs. This information enables the control of passive elements to achieve optimal signal reflection and manipulation. As narrow beams are required to overcome path loss, a high number of RIS elements is needed. This leads to a high control signaling overhead for controlling the RIS. The already challenging overhead is further affected by the following effects: In general, a cascaded channel between the base station, RIS, and UE must be estimated. However, obtaining accurate CSI of this channel is challenging due to potential near-field characteristics, mutual coupling, and beam squint effects, which can lead to errors in channel estimation and performance degradation. Furthermore, a recently published study on optimal RIS configuration in rich scattering environment indicates that the linear cascaded models are not correct [29, 30]. All named challenges add up, increasing the required overhead to optimal configure the RIS. To address the challenge of a tremendous overhead, passive beam training techniques employing predefined beam pattern codebooks were suggested [31, 32, 33, 34, 35, 36]. These approaches seek to exploit the inherent structure of the RIS-assisted channel to obtain accurate CSI required for optimal RIS manipulation but misses the ability for flexible deployments. Further research must investigate balancing the trade-off between RIS coverage and resource overhead while maintaining communication quality.
@@ II-C2 Regulatory and Multi-Operator Interference Issues
One of the major challenges for RIS deployment in multi-operator scenarios is managing interference among different network operators, ensuring fairness in resource allocation, and avoiding destructive interference. As RISs can potentially modify the propagation environment across multiple operators, the allocation of resources and the concurrent usage of multiple RISs deployed by different operators become critical concerns. The wideband properties and non-linear saturation effects of RISs introduce challenges related to unwanted interference and regulatory compliance [31]. When an RIS is deployed and controlled by one operator, it can cause unpredictable changes in the propagation environment for other operators in the nearby spectrum. Such alterations in the propagation environment can lead to issues such as additional reflections affecting cell coverage, neighbor-cell interference, and may disturb network planning. Furthermore, rapid variations in reflection patterns can disturb instantaneous spatial multiplexing and link adaptation for both reciprocity-based and downlink CSI-based scenarios. Research in this area is in its nascent stages, but preliminary work suggests the need for collaborative efforts among operators. Addressing these challenges necessitates the development of interference mitigation techniques and careful consideration of regulatory guidelines.
Possible solutions for mitigating this issue include the development of standardized protocols for RIS deployment and interference management, the enhancement of cooperation schemes like cooperative scheduling among network operators, and interference-aware resource allocation algorithms that minimize multi-operator interference. Further research must focus on identifying and quantifying the key factors influencing multi-operator interference and exploring advanced techniques for interference mitigation and management in RIS-assisted networks.
@@ II-C3 Competitive Technologies
[…]
The plot of $P$ for different frequencies is shown in Figure 3. The SISO case shows the highest transmit power $P$ for the considered distance $d_{1}$, increasing with increasing $d_{1}$. The relay shows a low and more or less constant value. The RIS case with RIS element number $N=25$ shows a slight improvement compared to the SISO case. The transmit power with RIS decreases when the RIS element number increases. Furthermore, if the RIS is close to the base station, the transmit power is low. Therefore, in this case, the RIS is competitive or even better than the relay where active gain of the signal is included. If the frequency increases, the power increases. In this case, relatively speaking, the relay needs less power even than the RIS with a higher element number. In summary, RIS is suitable for a position close to the BS (UE), where the advantage diminishes with increasing frequency.
(a) 6 GHz case.
(b) 27 GHz case.
Figure 3: Transmit power $P$ as a function of the distance $d_{1}$, which measures the distance between the source and the destination of the signal. The cases of the direct beam (SISO), a relay in between, and the RIS are considered. The RIS can be implemented with an element number $N$ of either $25$ or $250$ tiles per RIS.
The potential benefits of RISs must also be assessed compared to competitive technologies like decode-and-forward relays [37, 39] or network-controlled repeaters (NCRs) [40]. While these solutions may be more costly alternatives compared to RIS, they offer better controllability and potentially larger coverage, as RISs are more susceptible to multiplicative fading. The path loss of the BS – RIS or NCR – UE (BS–RIS/NCR–UE) link is a product of the individual link path losses. It can be compensated by amplification capabilities in the NCR case. In the case of RIS, it requires a close deployment of the RIS to either the BS or the UE to enable signal improvement or a very large RIS with an extensive element count (see also Figure 3). Besides the production costs, deployment issues, and maintenance costs, which all increase with size, a very large RIS leads to problems described in the earlier paragraph on channel estimation challenges. For a fair comparison between RIS and NCRs, all named advantages and disadvantages must be added up to a total cost of ownership per deployment. To evaluate the economic choice between RIS and NCR, network operators will have to consider the cost for coverage, which is a sum of the deployment costs per area. Because several cheaper RISs are required to cover the same area as one NCR, as stated earlier, the lower RIS deployment cost must level out this factor. The many aspects of the cost for coverage lead to a complex calculation. This evaluation is seen as a crucial aspect of future research. A more extensive comparison of RISs and NCRs is provided in [14].
@@ III Hardware Design
RISs are reflectors usually based on periodic structures that create a phase difference in the wave. The reconfigurability is obtained by modifying the phase difference between the different cells. The project 6G-LICRIS attempts to improve the state-of-the-art of RIS and investigate new hardware concepts. Some of the ideas being implemented within the scope of the project are briefly introduced. The proposed concepts follow a passive design, reflecting the outgoing RF wave and saving power compared to base station antennas, with the goal of enhancing coverage, spectral efficiency, and beam management while improving energy efficiency. Two different RIS models based on different approaches are proposed. The first model is based on resonant structures, while the second model is based on delay lines coupled to a patch antenna at the $28$ GHz band. Both RIS topologies are planned to be tested in indoor scenarios.
III-A Multi-resonance RIS
The first model consists of a multi-resonant RIS integrated into liquid crystal (LC) with wideband behavior and compatibility with ultra-thin layers. Since the switching time of the liquid crystal is directly proportional to the square of the thickness of the liquid crystal layer [41], it is important to maintain a thin LC layer to achieve fast switching and make the RIS suitable for changeable environments. This requirement may be problematic since very thin LC layers may create a mismatch between the impedance of the resonators and the free-space impedance. In [42], it is shown how multi-resonant structures can be used in combination with LC, but a LC layer with a thickness of around $0.05\lambda_{0}$ is used, and the project aimed at obtaining a working RIS with an LC thickness of around $0.002\lambda_{0}$, where $\lambda_{0}$ represents the free-space wavelength.
Some attempts were performed to achieve good performance with such a thin layer. For instance, in [43], the authors presented a RIS based on resonators compatible with thin LC layers, but the behavior is relatively narrowband. The presented topology consists of a layer of LC sandwiched between two layers of glass substrate. Different metallic resonant structures are printed onto the glass using photolithography processes. Therefore, they are placed directly on top and bottom of the liquid crystal, and their resonances are influenced by the polarization of the LC molecules. When multiple resonances are used, a wide bandwidth is expected. In this case, a bandwidth of around $6$ GHz with a central frequency of $28$ GHz is obtained by simulation. The proposed unit cell is compatible with a single linear polarization. Front and stack-up views of the RIS unit cell are shown in Figure 4.
(a) Stackup of the multi-resonance unit cell proposed.
(b) Schematic front view of the RIS, where the top resonant structures are shown.
Figure 4: Multi-resonance RIS design.
The expected performance of such a RIS was simulated using CST Studio [44], and the amplitude of the reflected wave as well as the phase for a single unit cell is shown in Figure 5, where the phase is normalized to the phase obtained for the minimum permittivity value. It can be observed that a wideband performance with a moderate amount of losses is obtained. The phase difference operates between approximately $24$ and $32$ GHz with a difference of around $320^{\circ}$ and a constant phase difference across the entire frequency range.
(a) Amplitude of the unit cell vs. frequency for different LC permittivity values.
(b) Phase of the unit cell vs frequency for different LC permittivity values.
Figure 5: Multi-resonance RIS simulation results.
III-B Delay-line based RIS
The second type of RIS consists of a coupled patch RIS. The working principle of this RIS topology is such that when the wave reaches the patch, it is coupled into the active phase shifter through two slots in the ground plane, where each slot works for a different polarization. In this topology, the radiating element is a patch that is coupled to a reflective LC phase shifter. Once the wave passes through the entire phase shifter, it is reflected at the end due to a termination in an open circuit, after which the wave is re-radiated. There have been some attempts to develop such a LC phase shifter-based RIS [45, 46], but due to space constraints, the lines are printed in a spiral shape, which does not allow for the use of dual polarization. To enable two polarizations, it is necessary to have the space to place two perpendicular phase shifters. Therefore, the lines should be miniaturized. The proposed method to achieve this is to design the phase shifter using miniaturization techniques, thereby increasing the phase shift per unit of length. The phase shifter may provide a variable phase shift ranging from $0^{\circ}$ to $360^{\circ}$. A bandwidth of around $3$ GHz is expected, and it is compatible with two orthogonal polarizations.
A preliminary view of the stack and a perspective view that allows seeing the patch over the two slots, which are coupled to the phase shifter, are shown in Figure 6. The performance of the unit cell can be seen in Figure 7. The amplitude and phase of the unit cell for different values of permittivity are shown, where the phase is normalized to the phase for the minimum permittivity value.
(a) Stackup view of the phaseshifter delay RIS.
(b) Top view of the unit cell based on a patch antenna over a glass substrate and a ground plane with two slots.
(c) View of the slots in combination with the LC-phase shifters topologies.
Figure 6: Delay-line based RIS design.
(a) Amplitude of the reflected wave at the unit cell for one polarization.
(b) Phase of the reflected wave at the unit cell for one polarization.
Figure 7: Delay-line based RIS simulation results.
@@ IV Deployment and Dimensioning Aspects
This section discusses the challenges related to RIS deployment and dimensioning. One of the main challenges in RIS deployment is related to the calculation of RIS placement in a cellular network, while RIS dimensioning involves calculating the number of reflecting elements necessary to achieve the required system targets. Central to deployment and dimensioning is the propagation model considered in the calculations. To this end, this section begins with a discussion on RIS propagation modeling, followed by deployment optimization, and RIS dimensioning.
@@ IV-A Propagation Modeling
In order to manage and control RISs from a system-level perspective, it is necessary to understand and ideally anticipate the effects of their configuration on intended changes in radio propagation as well as unintended effects, e.g., in-band and out-of-band interference. This requires fast and reliable RF propagation models, where a reasonable loss of accuracy may be acceptable.
Semi-empirical and/or geometric RF propagation models offer a good compromise. The measures that need to be taken into account are, first of all, the received power (see Figure 8). In a more sophisticated model, which is beyond the scope of the current work, the impact of the MIMO path through the RIS would also be of interest. Since RIS can be most efficiently used at higher frequencies beyond $6$ GHz, RF propagation in scenarios where RISs are used is typically limited by LOS between the transmitter and receiver. The intention of using RIS is to supplement the LOS paths with reflections off the RIS. While semi-empirical models are widely used in sub-$6$ GHz mobile radio scenarios (or historically sub-$2$ GHz), geometric models are less limited in their frequency application ranges, since they are based on deterministic approaches.
Figure 8: Path model of RF propagation components including an RIS working as a “directed mirror”. The blocks are named with “L” representing losses, “G” representing gains, and “P” representing power values. “Tx” stands for transmitting, and “Rx” for receiving. The loss of the RIS depends on the angle of arrival/transmission from the sender/receiver.
Semi-empirical models offer the advantage of relatively low computational complexity compared to other RF propagation modeling approaches, such as ray tracing and full-wave simulations, which can be computationally intensive and time-consuming. However, they also have limitations. These models are based on empirical measurements that may not always capture the full complexity of real-world environments, and they may not accurately predict the effects of new technologies and deployment scenarios that differ from the assumptions of the original model. Therefore, careful validation and calibration of semi-empirical models using real-world measurements and data are essential to ensure their accuracy and reliability.
@@
Geometric models are based on the concept that RF signals propagate in straight lines, and their behavior can be predicted by considering the geometry of the environment in which they propagate (see Figure 9, right-hand side). These models use parameters such as the distance between the transmitter and receiver, the height of the antennas, and the surrounding terrain to calculate the path loss of RF signals. For example, the two-beam ground reflection model accounts for reflections of RF signals from the ground by assuming that the signal travels along two paths: a direct path from the transmitter to the receiver and another path reflected from the ground to the receiver. The model considers the heights of the antennas and the distance between the transmitter and receiver to calculate the path loss. The two-beam ground reflection model is commonly used to predict RF propagation in urban environments where ground reflections can significantly affect path loss. A similar approach can be considered for RF predictions with RIS, assuming that the reflection point is from the RIS rather than the ground.
@@
More advanced geometric models can also consider additional factors such as the height and location of buildings, the density of obstructions, and the surrounding terrain. The Extended Hata Model [47] combines geometric principles with empirical data to account for the effects of both urban and suburban environments, including factors such as building density and road width. It can also account for the effects of obstacles, as is typically done in knife-edge models.
Knife-edge models are named for the sharp edge of the obstacle, which is considered the most significant factor affecting the diffraction of RF signals (see Figure 9, left-hand side). The shape, height, and distance of the obstacle from the transmitter and receiver are key parameters in these models. One of the most commonly used knife-edge models is the single knife-edge diffraction model, which assumes that the RF signal encounters a single sharp edge along its propagation path. The model calculates the path loss based on the height and distance of the obstacle, the height of the antennas, the frequency of the RF signal, and the wavelength of the signal.
For the integration of RIS into RF propagation models, diffraction is certainly of less interest than reflection. However, both modeling approaches could be combined through a combined parameter estimation of their parameters using methods of mathematical optimization, including machine learning/artificial intelligence approaches (see Figure 10). Such an approach strongly depends on the availability of suitable training data. For RIS, such data are hardly available. A combination of simulation and ray tracing may provide artificial training data, similar to the approach used by RadioUNet [48].
In any case, the RF characteristics of RIS need to be well understood for different bands, in various environments, and also under different environmental conditions. The proposed RF models need to be flexible in this respect, which outlines the need for simplified models. For control approaches, RF measurement data from the field must constantly feed the control loops, and the RF modeling on the RIS must also be adapted during operations. Therefore, an extensive control loop and measurement collection approach from connected devices within the network is required.